# Reactive Collision Avoidance for Safe Agile Navigation

### v1

@@ IV Experimental Setup and Implementation
Planning and Control. The NMPC framework employs $10$ shooting steps over a $1\leavevmode\nobreak\ $\mathrm{s}$$ prediction horizon, solved using sequential quadratic programming via acados[66], with a Gauss-Newton Hessian approximation and stabilization through Levenberg-Marquardt regularization. The control input vector $\mathbf{u}$ consists of the $4$ motor thrusts, governed by a dynamic model detailed in previous works[67, 68, 69]. The planned trajectory is generated using a min-jerk trajectory planner to ensure smooth transitions, while the desired control actions are derived through differential flatness principles. State and input weights are meticulously tuned for precise trajectory tracking. The NMPC operates efficiently with $10$ active CBF constraints—empirically validated as sufficient for our experiments—achieving average computation times of $2\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ per iteration.
@@ V Results and Performance Evaluation
Additionally, we analyze the effects of different polynomial orders for depth completion and the computational complexity for different camera resolutions and number of constraints.
@@ V-C Analysis: Polynomial Order for Depth Completion
We evaluate the impact of polynomial order on depth completion accuracy using NYUv2 [70] for indoor and KITTI [71] for outdoor data. The ground truth depth maps serve as sensed depth, and different polynomial orders are fitted to the monocular depth estimates and ground truth to determine the optimal degree for our experiments. Key metrics include Root Mean Square Error (RMSE), which highlights larger depth discrepancies, and Absolute Relative Error (AbsRel), reflecting proportional accuracy.
The results in Table I indicate that the quadratic model achieves the lowest errors on both datasets, outperforming the linear model, which fails to capture non-linear variations, and the cubic model, which tends to overfit.
@@ V-D Analysis: Computational Complexity
Table I: Analysis on polynomial order for depth completion. Performance of different polynomial orders (linear, quadratic, and cubic) used for depth completion on the NYUv2 and KITTI datasets. The quadratic model consistently outperforms the others, achieving the best accuracy across various metrics and datasets.
@@ VI Conclusion and Future Directions
This paper presented a new obstacle avoidance framework that integrates perception constraints directly into an NMPC, enabling agile and safe navigation for autonomous robots. By jointly addressing perception, planning, and control, the proposed approach effectively navigated complex environments in real-time, as demonstrated through extensive indoor and outdoor experiments. The system’s ability to dynamically adjust control barrier functions ensured robust performance without reliance on specific environmental knowledge, making it suitable for various applications, including target tracking and autonomous exploration.
While our current work primarily focused on navigating cluttered environments with significant noise and environmental disturbances, future work will extend the system to handle dynamic obstacles by integrating obstacle kinematics prediction into the NMPC. We also plan to transition to vision-based navigation, allowing the quadrotor to rely solely on camera inputs, which will simplify the hardware and expand its applicability to compact and lightweight scenarios.

### v3

@@ IV Experimental Setup and Implementation
Planning and Control. The NMPC framework uses $10$ shooting steps over a $1\leavevmode\nobreak\ $\mathrm{s}$$ prediction horizon, solved via sequential quadratic programming with acados [66], leveraging a Gauss-Newton Hessian approximation and Levenberg-Marquardt regularization. The control input $\mathbf{u}$ includes four motor thrusts, governed by a dynamic model from previous works [67, 68, 69]. A min-jerk trajectory planner ensures smooth transitions, while differential flatness principles derive the desired control actions. Carefully tuned state and input weights enable precise trajectory tracking. The NMPC runs efficiently with $10$ active CBF constraints, validated experimentally, achieving an average computation time of $2\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ per iteration.
@@ V Results and Performance Evaluation
Additionally, we analyze the impact of polynomial order on depth completion and the computational complexity across camera resolutions and constraint numbers.
@@ V-C Analysis: Polynomial Order for Depth Completion
We analyze the impact of polynomial order on depth completion accuracy using NYUv2 [70] (indoor) and KITTI [71] (outdoor). Ground truth depth maps serve as the reference, and polynomial functions of different orders are fitted to monocular depth estimates. Key metrics include Root Mean Square Error (RMSE), which emphasizes large depth discrepancies, and Absolute Relative Error (AbsRel), which captures proportional accuracy.
As shown in Table I, the quadratic model achieves the lowest errors on NYUv2, while the linear model performs better on KITTI in terms of RMSE and AbsRel, likely due to the nature of outdoor depth variations and sparser LiDAR-based ground truth. The cubic model tends to overfit in both cases. Additionally, KITTI exhibits higher RMSE and AbsRel values due to its greater depth range and sparser labels compared to the denser RGBD data in NYUv2.
@@ V-D Analysis: Computational Complexity
Table I: Analysis on polynomial order for depth completion. Performance of different polynomial orders (linear, quadratic, and cubic) used for depth completion on different datasets. The quadratic model achieves the best accuracy on NYUv2, while the linear model performs slightly better on KITTI in terms of RMSE and AbsRel.
@@ VI Conclusion and Future Directions
This paper presented a new obstacle avoidance framework that integrates perception constraints directly into an NMPC, enabling agile and safe navigation for autonomous robots. By tightly addressing perception, planning, and control, the proposed approach effectively navigated complex environments in real-time, as demonstrated through extensive indoor and outdoor experiments. The system’s ability to dynamically adjust control barrier functions ensured robust performance without reliance on specific environmental knowledge, making it suitable for various applications, including target tracking and autonomous exploration.
Future work will extend the system to handle dynamic obstacles by incorporating obstacle kinematics prediction into NMPC. Additionally, transitioning to vision-based navigation will allow the quadrotor to rely solely on camera inputs, reducing hardware complexity and enhancing applicability to compact scenarios.